OpenStudy (anonymous):

An urn contains six red balls, five white balls, and four black balls. Four balls are drawn from the urn at random without replacement. For each red ball drawn, you win $6, and for each black ball drawn, you lose $9. Let X represent your net winnings.

hartnn (hartnn):

what do u need to find ? expected profit ?

OpenStudy (anonymous):

That's correct. Sorry, I should have put that in there.

hartnn (hartnn):

Can you find the probability of selecting a red ball from the urn ?

hartnn (hartnn):

I also need some more info. how many times in total, you will be drawing the ball ?

hartnn (hartnn):

ok, 4 times

OpenStudy (anonymous):

Only once with no replacement I believe.

hartnn (hartnn):

Find total number of balls in the urn . P(red ball) = total number of red balls / total number of balls in the urn = ...

OpenStudy (anonymous):

Oh, you meant how many times a single ball is drawn. Then yes, you're correct. 4 times.

OpenStudy (anonymous):

P(6)=6/15=.4

hartnn (hartnn):

good, so you know how to calculate probabilities. ok, in first draw, there are 15 balls, you can pick a red OR a black ball. so your expected profit in 1st draw will be P(red)*6 - P(black)*9 The second term is subtracted as you lose $9 if you draw a black ball does that make sense ?

OpenStudy (anonymous):

P(black)=4/15=.3 correct?

hartnn (hartnn):

yes bu note that what i stated is true only for one(first) draw

hartnn (hartnn):

Get the probability of drawing 4 red balls in 4 draws of balls from the urn without replacement

OpenStudy (kropot72):

There are 15 possible combinations of red and black, as follows: R B 4 0 3 0 2 0 1 0 0 0 3 1 2 1 1 1 0 1 2 2 1 2 0 2 1 3 0 3 0 4 For starters, the probability of each needs to be found.

OpenStudy (anonymous):

Wait, isn't this where I would do something like 4/15 * 3/14 * 2/13 * 1/12 ?

hartnn (hartnn):

As you know 0.4 is just the probability of drawing one red ball from the urn in one draw. Here's what you need to do, Take a row from the above table krotop posted, say 4,0 'Get the probability of drawing 4 red balls in 4 draws of balls from the urn without replacement' which is what i asked. and your reply '4/15 * 3/14 * 2/13 * 1/12' for this is correct :) Repeat this for all the rows mentioned in that table

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is going to take a minute. LOL. I know there's a quicker way to do this using the nCr function on my calculator, but I've been doing this stuff all day and I'm starting to confuse myself. lol

hartnn (hartnn):

Once you get the probabilities, multiply it with your profit! like for the first row, your profit will be 4*6 =$24 multiply 24 with the probability of 1st row of you take the row with 1,3 1 red and 3 balck = +6 - 3*9 = -21 multiply -21 with the probability you calculated for 1,3 repeat for all rows, and add all the profits

hartnn (hartnn):

do you have some idea on what you need to do /

OpenStudy (anonymous):

Oh ok. Gotcha. Just doing the math now.

OpenStudy (anonymous):

Just to make sure I'm doing this right before I get too far, the answer for the first row would be 83865.6?

OpenStudy (kropot72):

For the first row, the probability of drawing 4 red is 0.01099. Then you multiply that value of probability by $24.

OpenStudy (anonymous):

Ok, I was doing it wrong. Didn't get 0.01099

OpenStudy (kropot72):

The probability of drawing 4 red is given by: \[\large P(4\ red)=\frac{6C4}{15C4}\]

OpenStudy (anonymous):

Ok, now I got 0.01099.

OpenStudy (anonymous):

The same would be for the second row, but instead I would use P(3red)=6C3/15C4 correct?

OpenStudy (kropot72):

The probability of drawing 3 red, 1 white and 0 black is given by: \[\large P(3\ red,\ 1white,\ 0\ black)=\frac{6C3 \times 5C1}{15C4}\]

OpenStudy (anonymous):

Yep, gotcha. Thanks.

OpenStudy (kropot72):

yw

OpenStudy (anonymous):

the very last row is giving me an answer with \[\epsilon ^{-4}\]

OpenStudy (anonymous):

.0282 is what I got as my final answer. Hoping that was right.

OpenStudy (kropot72):

The probability of 4 black and 0 red is given by: \[\large P(4\ black)=\frac{1}{15C4}=0.000733\]